# Bio-Inspired Classification: Combining Information Theory and Spiking Neural Networks - Influence of the Learning Rules

### v1

@@ IV Results and Discussion
TABLE IV: Influence of hybrid learning algorithm on SNN.
Type Subtype Datasets Bio-inspired Epochs $n$ Time [s] Accuracy [%] MSE MAE $R^{2}$ score
Hybrid ANN-SNN Bernoulli No 10 64 21.41 99.50 0.0050 0.0050 0.9800
Markov No 10 32 37.25 92.58 0.0742 0.0724 0.7031
Poisson No 10 32 115.41 94.50 0.0550 0.0550 0.7799
Reward Bernoulli No 10 32 4.3 99.50 0.0050 0.0050 0.9800
Markov No 10 64 36.32 90.35 0.0965 0.0965 0.9948
Poisson No 10 64 14.83 91.00 0.0900 0.0900 0.6399
BAL Bernoulli Yes 10 32 1.56 97.00 0.0300 0.0300 0.8800
Markov Yes 10 128 46.37 95.38 0.0462 0.0462 0.8153
Poisson Yes 10 128 95.98 94.50 0.0550 0.0550 0.7799
V Conclusions
This comprehensive benchmarking study sheds light on the capabilities, trade-offs, and biological plausibility of various learning strategies in Spiking Neural Networks. It reveals that no single algorithm universally outperforms others across all metrics, data types, or computational conditions.
Higher input correlations, as observed for Poisson-distributed data, tend to reduce classification accuracy, emphasizing the inherent difficulty of extracting informative features from less structured signals. In contrast, more structured input types such as Bernoulli and Markov sequences facilitate greater learning stability and improved performance.
Hebbian learning remains computationally efficient and accurate for structured data, but struggles with Poisson-distributed inputs. STDP and SDSP, grounded in spike timing, provide biologically plausible yet effective mechanisms for learning from both structured and moderately unstructured inputs. Algorithms like BP and STBP are unmatched in precision but are computationally expensive, particularly for inputs like Poisson sequences. Their reliance on full backpropagation limits real-time neuromorphic applicability, despite theoretical optimality. ANN-SNN conversion trough a LZC and reward-based STDP offer scalable and flexible solutions. BAL, in particular, balances biological realism and computational efficiency, emerging as a promising method for future biologically plausible SNNs. Poisson inputs offer a realistic benchmark but challenge traditional learning rules. Hybrid and bio-inspired algorithms like BAL and SDSP demonstrate strong adaptability to these conditions, offering a viable path toward more robust neuromorphic learning systems.
Flexible hybrid learning architectures that dynamically adapt learning strategies in response to input characteristics hold promise for enhancing the performance of spiking neural networks under diverse operational conditions. Advancing our understanding of spike contributions in non-differentiable learning contexts remains a key limitation. Integrating event-driven learning mechanisms with biologically inspired regularization may offer promising directions for enhancing optimization efficiency in SNNs. Moreover, real-time evaluation of the trade-offs between accuracy and computational efficiency is essential for effective integration of SNNs on neuromorphic hardware, particularly in edge and low-power computing environments. To approach functional parity with biological intelligence, future learning models must better emulate the noise robustness and generalization capabilities of the brain, in particular under stochastic, Poisson-like input regimes.

### v2

@@ 5 Results and Discussion
Table 6: Wilcoxon signed-rank test for synthetic stochastic datasets (Bernoulli, Markov, Poisson). All comparisons yield $p>0.05$, indicating no statistically significant differences between learning rules.
Dataset Comparison $p$ (LZC acc.)
Bernoulli All algorithm pairs >0.05>0.05
Markov All algorithm pairs >0.05>0.05
Poisson All algorithm pairs >0.05>0.05
Table 7: Wilcoxon signed-rank test for pairwise comparisons between learning rules (Tempotron, SpikeProp, BAL) on neuromorphic datasets. We report $p$-values for LZC-based accuracy and raw SNN accuracy. Bold values indicate statistically significant differences at $\alpha=0.05$.
Dataset Comparison $p$ (LZC acc.) $p$ (SNN acc.)
MNIST01 Tempotron vs SpikeProp 0.32 0.32
MNIST01 Tempotron vs BAL n/a n/a
MNIST01 SpikeProp vs BAL 0.32 0.32
FashionMNIST01 Tempotron vs SpikeProp 0.11 0.07
FashionMNIST01 Tempotron vs BAL 0.32 0.18
FashionMNIST01 SpikeProp vs BAL 0.10 0.07
N-MNIST01 Tempotron vs SpikeProp 0.89 0.69
N-MNIST01 Tempotron vs BAL 0.72 0.028
N-MNIST01 SpikeProp vs BAL 0.83 0.028
Gesture01 Tempotron vs SpikeProp 0.18 0.59
Gesture01 Tempotron vs BAL 0.32 0.47
Gesture01 SpikeProp vs BAL 1.00 0.27
SHD01 All pairs n/a11footnotemark: 1 n/a22footnotemark: 2
Tables 6 and 7 summarize the statistical reliability of LZC-based classification using Wilcoxon signed-rank tests across three SNN training algorithms (Tempotron, SpikeProp, BAL) and multiple data sources. For synthetic stochastic datasets (Bernoulli, Markov, Poisson), all pairwise comparisons yield $p>0.05$, indicating no statistically significant differences between learning rules. A similar trend is observed for neuromorphic datasets, where most comparisons remain statistically indistinguishable, with only isolated differences appearing in raw SNN accuracy rather than in LZC-based decisions. These results support the observation that LZC representations produced by different learning rules remain statistically comparable across datasets. Non-significant Wilcoxon results confirm that differences between learning rules are not statistically supported at the level of LZC-based decisions.
Table 8: Influence of hybrid learning algorithms on SNN performance.
Type Subtype Dataset Bio- Ep. $n$ Time [s] Acc. [%] MSE $R^{2}$
Hybrid ANN–SNN Bernoulli No 10 64 $21.41$ $99.50$ $0.0050$ $0.9800$
Markov No 10 32 $37.25$ $92.58$ $0.0742$ $0.7031$
Poisson No 10 32 $115.41$ $94.50$ $0.0550$ $0.7799$
Reward Bernoulli No 10 32 $4.30$ $99.50$ $0.0050$ $0.9800$
Markov No 10 64 $36.32$ $90.35$ $0.0965$ $0.9948$
Poisson No 10 64 $14.83$ $91.00$ $0.0900$ $0.6399$
BAL Bernoulli Yes 10 32 $1.56$ $97.00$ $0.0300$ $0.8800$
Markov Yes 10 128 $46.37$ $95.38$ $0.0462$ $0.8153$
Poisson Yes 10 128 $95.98$ $94.50$ $0.0550$ $0.7799$
Table 9 provides computational metrics including FLOPs per forward pass, runtime, latency per processed sample, and energy estimates. Energy values are analytical estimates derived from runtime and an assumed average power draw of the measurement platform; they should be interpreted as coarse proxies rather than hardware-measured neuromorphic energy. These results demonstrate that all three SNN learning rules operate in a low-FLOPs regime, with per-sample latency below 0.25 ms for synthetic data and below 0.1 ms for MNIST after optimization, supporting real-time feasibility. It is important to emphasize that Lempel–Ziv complexity is a single-scalar descriptor of temporal structure rather than a discriminative classifier optimized for decision boundaries. When used with a fixed threshold, LZC-based decisions can become unstable under class imbalance or when class-conditional complexity distributions overlap, which may reduce the F1-score even if overall accuracy remains high. In this paper, we therefore interpret LZC primarily as a decision-relevant representation that exposes how learning rules shape temporal information, while standard metrics (accuracy, MSE, $R^{2}$) quantify separability under the adopted protocol.
Table 9: Computational efficiency metrics: runtime, latency, FLOPs, and estimated energy per sample.
Dataset Algorithm Time per run [s] Latency [ms/sample] FLOPs [G]/forward Energy [mJ/sample]
(mean) (mean $\pm$ std) (mean) (mean)
Bernoulli Tempotron $46.67$ $233.34$ p m 4.17 0.000 2950.000\,295 0.002 9490.002\,949
SpikeProp $46.63$ $233.16$ p m 3.74 0.000 2950.000\,295 0.002 9490.002\,949
BAL $23.16$ $115.81$ p m 42.95 0.000 2950.000\,295 0.002 9490.002\,949
Markov Tempotron $46.77$ $233.87$ p m 7.12 0.000 2950.000\,295 0.002 9490.002\,949
SpikeProp $47.31$ $236.54$ p m 6.95 0.000 2950.000\,295 0.002 9490.002\,949
BAL $30.75$ $153.77$ p m 65.54 0.000 2950.000\,295 0.002 9490.002\,949
Poisson Tempotron $45.15$ $225.75$ p m 3.76 0.000 2950.000\,295 0.002 9490.002\,949
SpikeProp $45.24$ $226.20$ p m 2.48 0.000 2950.000\,295 0.002 9490.002\,949
BAL $28.93$ $144.65$ p m 70.30 0.000 2950.000\,295 0.002 9490.002\,949
MNIST 0/1 Tempotron $208.96$ $70.69$ p m 9.00 0.000 2330.000\,233 0.002 3350.002\,335
SpikeProp $105.49$ $35.69$ p m 21.89 0.000 2330.000\,233 0.002 3350.002\,335
BAL $71.39$ $24.15$ p m 2.91 0.000 2330.000\,233 0.002 3350.002\,335
Our experiments reveal that the choice of learning algorithm must be carefully aligned with both the structure of the input data and application-specific computational constraints. Classic backpropagation and surrogate-gradient methods (e.g., STBP) achieve superior accuracy, particularly for structured signals. However, their computational demands are prohibitively high for real-time or resource-constrained deployments. On the other hand, bio-inspired learning algorithms such as tempotron, SDSP, and BAL offer a more favorable trade-off, delivering robust classification with substantially lower computational costs. BAL, in particular, emerges as a promising candidate, balancing biological plausibility with operational efficiency.
Analysis across input types shows that classification accuracy is strongly influenced by the temporal correlation structure of the signals. Poisson-distributed inputs, despite being memoryless and weakly correlated, pose the greatest challenge for learning algorithms due to their high temporal variability and lack of exploitable structure. In contrast, Markovian sequences, with moderate temporal dependencies, offer a balanced scenario, while Bernoulli sequences, being memoryless and maximally entropic per symbol, provide a broad baseline for learning without temporal dependencies. These findings highlight the importance of aligning learning strategies with the statistical properties of input, especially in the design of SNNs for real-world data processing.
In general, achieving brain-like robustness under noisy stochastic input regimes, particularly those modeled by Poisson processes, will require continued refinement of adaptive, efficient, and biologically plausible learning strategies. The LZC-SNN framework proposed here contributes to this goal by offering a scalable, interpretable, and resource-efficient alternative to conventional classification pipelines in spiking neural systems.
While our previous study examined the role of neuron models, the present work demonstrates that learning paradigms exert an equally strong and often dominant influence on the temporal organization of spiking activity, as revealed by complexity-based analysis.
5.1 Limitations of fixed-threshold LZC decisions
Lempel-Ziv complexity is a single-scalar summary of temporal novelty and is not optimized to directly maximize discriminative margins. When class priors are imbalanced or when class-conditional LZC distributions overlap, a fixed decision threshold may collapse into predicting the majority class. This leads to deceptively high classification accuracy while driving the F1-score toward zero, indicating poor minority-class recognition.
For this reason, LZC is interpreted in this work primarily as a decision-relevant temporal representation that exposes how learning rules reorganize spike-train structure, rather than as a strong stand-alone classifier. Classification metrics are therefore used as downstream probes of separability under a specified protocol, and class-sensitive measures are reported to explicitly flag regimes in which scalar-complexity decisions become unreliable.
6 Conclusions
This study set out to examine the role of learning algorithms in shaping the classification performance and computational efficiency of Spiking Neural Networks. By systematically comparing classical, bioinspired, and hybrid training approaches across datasets with varying temporal structures Bernoulli, Markov, and Poisson processes we were able to characterize how algorithmic choices interact with data statistics to determine both accuracy and runtime.
Our results provide several insights. First, gradient-based methods such as backpropagation and surrogate-gradient learning remain unrivalled in terms of predictive precision, consistently achieving near-perfect accuracy and negligible error values. However, their computational costs were prohibitively high, in some cases exceeding 48,000 seconds (measured over 10 epochs on a standard desktop machine with Intel® Core™ i7-14700F @ 2.10 GHz), which makes them impractical for real-time or embedded neuromorphic deployments. Second, bioinspired algorithms such as Hebbian learning, STDP, SDSP, tempotron, and SpikProp offered substantially lower computational costs, often two orders of magnitude faster, while still maintaining competitive accuracy for structured sequences. Their limitations became apparent in the context of Poisson-distributed data, where stochastic variability hindered generalization and accuracy degraded. Third, hybrid schemes including ANN-SNN conversion and BAL were shown to mitigate some of these weaknesses by leveraging biologically plausible mechanisms that adapt more flexibly to irregular temporal inputs.
The methodological novelty of this work lies in the use of Lempel-Ziv complexity as a decision-relevant temporal descriptor embedded within an SNN framework. Unlike prior studies where LZC served primarily as a descriptive post hoc metric, here it is used to analyze how different learning rules shape temporal representations and influence classification-related outcomes, rather than as a standalone discriminative classifier. This integration of information-theoretic complexity analysis with biologically plausible spiking computation advances the field in two important ways: it improves interpretability by linking decision boundaries to structural novelty in the input, and it enhances efficiency by reducing reliance on gradient-based training. We believe this conceptual bridge between complexity theory and neuromorphic computing can serve as a useful template for future research in pattern recognition with SNNs.
At the same time, several limitations must be acknowledged. The experiments were conducted on synthetic datasets, which, while useful for systematically controlling temporal correlations, may not capture the full irregularities of real-world neural or sensory data. Moreover, although LZC-SNN achieved competitive accuracy, the method requires further validation on larger-scale problems and with hardware implementations to fully assess scalability. Finally, while we report coarse energy estimates, hardware-measured energy on neuromorphic processors remains out of scope and will be addressed in future work.
Despite these limitations, the results carry implications for both researchers and practitioners. For researchers, the comparative framework presented here provides a benchmark for assessing novel SNN learning rules under varying temporal conditions. For practitioners designing neuromorphic systems, the findings clarify trade-offs between accuracy and efficiency and highlight which algorithms are best suited for structured versus stochastic inputs.
The main strength of our approach lies in its interpretability efficiency compared to conventional decoding strategies. At the same time, its accuracy is sensitive to the statistical properties of the input and to the choice of learning rules, which we identified as a limiting factor for real-world deployment. These insights can guide researchers in selecting appropriate training strategies depending on task constraints.
Future work should proceed along several directions. Extending the evaluation to real-world datasets such as electrophysiological recordings, speech signals, or event-based vision would test the robustness of the observed trends. Hardware level experiments are needed to quantify energy savings and latency reductions in neuromorphic processors. Methodologically, there is an opportunity to design adaptive or hybrid algorithms that dynamically switch between gradient-based and bioinspired updates depending on the input statistics, thereby combining accuracy with efficiency. Finally, further exploration of information-theoretic measures beyond LZC, such as entropy rate or predictive information, may provide complementary tools for interpretable spiking computation.
In summary, this work highlights the decisive role of learning algorithms in determining the balance between accuracy, efficiency, and biological plausibility in SNNs. By embedding complexity-theoretic principles into spiking architectures, we take a step toward interpretable and resource-efficient pattern recognition systems that more closely align with the constraints of real-world neuromorphic applications.
7 Practical Implications
The findings suggest that classical gradient-based algorithms should be reserved for offline or high-resource scenarios requiring maximal accuracy, whereas bioinspired and hybrid methods provide a more favorable balance for real-time and resource constrained neuromorphic applications. The proposed LZC-SNN framework, by combining spiking dynamics with information-theoretic analysis, offers a scalable and interpretable approach to temporal pattern recognition, with potential applications in brain-computer interfaces, biomedical signal analysis, and event-driven sensory processing.